# Binary Event-Driven Spiking Transformer

### v1

@@ Abstract
Transformer-based Spiking Neural Networks (SNNs) introduce a novel event-driven self-attention paradigm that combines the high performance of Transformers with the energy efficiency of SNNs. However, the larger model size and increased computational demands of the Transformer structure limit their practicality in resource-constrained scenarios. In this paper, we integrate binarization techniques into Transformer-based SNNs and propose the Binary Event-Driven Spiking Transformer, i.e. BESTformer. The proposed BESTformer can significantly reduce storage and computational demands by representing weights and attention maps with a mere 1-bit. However, BESTformer suffers from a severe performance drop from its full-precision counterpart due to the limited representation capability of binarization. To address this issue, we propose a Coupled Information Enhancement (CIE) method, which consists of a reversible framework and information enhancement distillation. By maximizing the mutual information between the binary model and its full-precision counterpart, the CIE method effectively mitigates the performance degradation of the BESTformer. Extensive experiments on static and neuromorphic datasets demonstrate that our method achieves superior performance to other binary SNNs, showcasing its potential as a compact yet high-performance model for resource-limited edge devices.
@@ 1 Introduction
Spiking Neural Networks (SNNs) have attracted significant attention as third-generation artificial neural networks, known for their high biological plausibility and low power consumption Maass (1997). The spiking neuron utilizes binary spikes as the fundamental units for information transmission and works in a sparse spike-driven manner Zhang et al. (2021). This sparse synaptic transmission in SNNs simplifies multiply-accumulate (MAC) operations into accumulate (AC) operations, thereby significantly enhancing computational efficiency Li et al. (2023); Xu et al. (2024a). Furthermore, the energy-efficient feature of SNNs has driven the development of neuromorphic hardware, such as SpiNNaker Painkras et al. (2013), TrueNorth Akopyan et al. (2015), Loihi Davies et al. (2018), and Tianjic Pei et al. (2019). However, despite the notable energy efficiency of SNNs, their performance in complex tasks still requires improvement.
In recent years, there have been several studies integrating Transformers into SNNs, leading to a series of high-performance models, such as Spikformer v1 and v2 Zhou et al. (2023b, 2024), Spkingformer Zhou et al. (2023a), Spike-Driven Transformer v1 and v2 Yao et al. (2024a, b), and SpikingResformer Shi et al. (2024). Compared to convolutional architectures in SNNs, these Transformer-based models have demonstrated significant performance improvements Zhang et al. (2022). However, their advancements typically rely on large model size, which is accompanied by substantial memory storage and computational overhead, limiting their deployment on resource-constrained edge devices. Therefore, there is an urgent need for a compact yet high-performance Transformer-based SNN.
Figure 1: Accuracy vs. NS-ACE & Model Size. Our method achieves superior computational and storage efficiency while outperforming other quantized SNNs on ImageNet. Neuromorphic Synaptic Arithmetic Computation Effort (NS-ACE) assesses SNN resource use in neuromorphic computing environments.
Quantization is a highly effective method for compressing large-scale models, which reduces model parameters from 32-bit to a low bit-width representation Yang et al. (2019). As an extreme form of quantization, binarization maximizes model size compression and accelerates computational speed by employing bitwise operations Qin et al. (2020a). Therefore, incorporating binarization with Transformer-based SNNs is promising for achieving an efficient and high-performance model. It is worth noting that current research on binarization in SNN domains is primarily focused on convolutional structures, while Transformer-based structures remain unexplored Chowdhury et al. (2021); Yin et al. (2024); Wei et al. (2024).
@@ 2.1 Transformer-based SNNs
In recent years, there have been several studies integrating Transformer architectures into SNNs, leading to a series of high-performance SNN models. Spikeformer Li et al. (2022b) is the first to integrate the Transformer architecture with SNNs, however, it retains numerous floating-point operations, making it unsuitable for neuromorphic computation. Spikformer Zhou et al. (2023b) introduces the Spiking Self Attention (SSA) mechanism, which enhances both energy efficiency and performance of Transformer-based SNNs. Based on this, Spikingformer Zhou et al. (2023a) modifies the residual connection within it to achieve a purely spike-driven Vision Transformer, further enhancing model’s efficiency. Spike-driven Transformer Yao et al. (2024b) proposes a spike-driven self-attention mechanism with linear complexity, significantly reducing energy consumption. Then, to ensure versatility and high performance across various vision tasks, Yao et al. (2024a) expand the original architecture into Meta-SpikeFormer, also known as Spike-driven Transformer v2. Moreover, SpikingResformer Shi et al. (2024) combines the ResNet architecture with a novel spiking self-attention mechanism, further improving performance and energy efficiency. Despite much progress, these models are limited by substantial memory and computational overheads, underscoring the need for further compression to reach their full potential.
@@ 2.2 Quantization techniques in SNNs
Various approaches have been proposed to quantize SNNs to low-bits. Deng et al. (2021) employ spatiotemporal backpropagation (STBP) to directly train quantized SNNs and introduce the alternating direction method of multipliers (ADMM) to solve the performance degradation caused by quantization. Then, to further enhance the performance, Yoo and Jeong (2023) use constrained backpropagation (CBP) with the Lagrangian function as an objective function to quantize SNNs in training. As an extreme form of quantization, binarization has also been widely studied. Qiao et al. (2021) present a weight-binarized SNN to efficiently process event-based data, addressing the training demand of neuromorphic hardware for event data. Moreover, Pei et al. (2023) propose the accuracy loss estimator and binary weight optimization to achieve ultra-low latency adaptive local binary SNNs, which reduce memory storage by over 20% while still maintaining high recognition accuracy. Recently, Wei et al. (2024) introduce a quantized SNN (Q-SNN) that reduces both weight and membrane potential representation, they also propose a weight-spike dual regulation (WS-DR) method to enhance the performance of Q-SNN. Despite the effectiveness of these binarization methods in SNNs, they mainly focus on spiking convolutional architectures, while Transformer-based SNNs have not been explored.
@@ 3 Method
Table 1: A simple example of binarizing Attn with $\operatorname{boolean}$ function, HR-LIF, and SR-LIF. The number of split patches $N$, time step $T$, threshold $V_{th}$, and time constant $\tau$ are set to 1, 4, 1, 0.5, respectively.
@@ 3.1.1 Weight binarization
where $\tau$ is time constant factor, $\text{U}_{l}[t]$ is the membrane potential of neurons in layer $l$ at time $t-1$, $\tilde{\text{U}}_{l}[t]$ is its intermediate representation, and $\text{X}_{l}[t]\in\mathbb{R}^{C\times H\times W}$ is the input current. $\text{X}_{l}[t]$ is integrated by presynaptic neurons, described as:
@@
Neurons reset its membrane potential after emitting a spike. Typically, we set reset potential to 0:
@@
where $\otimes$ is efficient bitwise operations, theoretically offering a 32×32$\times$32 × memory saving and speedup compared with 32-bit operations Hubara et al. (2016); Rastegari et al. (2016a).
Figure 2: The ‘LIF-B-Conv-BN’ structure of BESTformer and representation capability of variables in the structure. Value set indecates the collection of all values present in a variable. Set size indicates the size of a value set.
@@ 3.1.2 Attention binarization
[…]
@@ 3.2 Challenge analysis
Previous research indicates that the poor performance of binarized neural networks is attributable to their low representational capability Liu et al. (2018); Guo et al. (2022, 2024). In our research, we experimentally analyzed the average representation capability of the full-precision network and BESTformer on ImageNet-1k, respectively. We use the value set size of feature maps as the measure of representation capability, which refers to the maximum number of distinct values in a feature map. The specific results are shown in the Figure 2.
@@ 3.3 Coupled information enhancement BESTformer
Due to the constrained information representation capability of the binarized model, the ideal binary model should aim to retain the information representation of their full-precision counterparts as much as possible, thus the mutual information between the binarized and full-precision models’ representations should be maximized Qin et al. (2022); Li et al. (2022a); Xu et al. (2023, 2024b). Therefore, we propose a knowledge distillation framework, optimizing the mutual information $\mathcal{I}$ between the student model $\mathcal{S}$ and the full-precision teacher model $\mathcal{T}$, which can be formalized as:
@@
$\mathcal{I}(\text{X}_{n}^{\mathcal{S}};\text{X}_{m}^{\mathcal{T}})=\mathcal{H}(\text{X}_{n}^{\mathcal{S}})-\mathcal{H}(\text{X}_{n}^{\mathcal{S}}|\text{X}_{m}^{\mathcal{T}}).$ (11)
We employed coupled information enhancement methods to optimize the objective: (1) maximizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$ to its upper bound by reversible framework, and (2) minimizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}}|\text{X}_{m}^{\mathcal{T}})$ by information-enhanced distillation. The overall architecture diagram of applying the CIE method to BESTformer is shown in Figure 3.
@@ 3.3.1 Reversible framework
The information entropy of the model’s feature maps exhibits a decreasing trend as the number of layers in the BESTformer increases. This can be shown explicitly by the inequality in proposition 1 and the detailed proof of proposition 1 is given in Appendix 6.1.
@@
Proposition 1 shows that the information retained by model tends to decrease as the neural network deepens, which is contrary to our objective of maximizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$. To address this issue, we utilize a reversible forward mapping in the encoder, as the reversible connection shown in Figure 3. We define forward mapping $\Phi_{l}(\text{X}_{l-1}^{\mathcal{S},0},\text{X}_{l-1}^{\mathcal{S},1})=(\text{X}_{l}^{\mathcal{S},0},\text{X}_{l}^{\mathcal{S},1})$ as:
@@
Inverse mapping does not directly engage in network’s operations, but it indirectly confirms that the information entropy within the reversible framework remains constant across all encoder blocks. Therefore, $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$ reaches its upper bound $\mathcal{H}(\text{X}_{0}^{\mathcal{S}})$ as demonstrated by proposition 2, which aligns with the objective of maximizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$. The detailed proof of proposition 2 is given in Appendix 6.2.
@@
Figure 4: Illustration of the forward and inverse process of the proposed reversible framework. The inverse process indicates that the inputs can be reconstructed from the outputs, i.e. this framework is reversible and no information lost.
@@ 3.3.2 Information enhanced distillation
This dual-head design offers a significant advantage over traditional distillation methods by decoupling the gradient backpropagation for distillation and classification at the head level. This decoupling reduces mutual interference between the two gradients during backpropagation, facilitating more effective parameter updates. It is worth noting that $\text{X}_{n}^{\mathcal{S},1}$, being the output of a deeper network compared to $\text{X}_{n}^{\mathcal{S},0}$, is more susceptible to overfitting when used for classification. Therefore, we strategically feed $\text{X}_{n}^{\mathcal{S},1}$ into the distillation head. This approach leverages the teacher model’s output, which encapsulates ‘error information’, thereby alleviating the overfitting problem and enhancing the model’s generalization capabilities.
@@ 4 Experiments
In this section, we first assess the classification performance of the proposed BESTformer with the CIE method on small-scale datasets including CIFAR Krizhevsky et al. (2009), CIFAR10-DVS Li et al. (2017). Following this, we evaluate the method’s performance on large-scale image dataset, ImageNet-1K Deng et al. (2009) to verify the scalability of our approach. Finally, we perform a series of ablation studies to validate the effectiveness of our method. The implementation details is provided in Appendix 6.3.
@@ 4.1.2 Results on ImageNet-1k classification
Compared to full-precision models, our method significantly reduces model size and computational efficiency. According to Table 3, our 1-bit Bestformer-8-512 model is 5.57MB, approximately 10 times smaller than its full-precision counterpart (56.61MB). In terms of computational efficiency, our 1-bit Spikingformer-8-512 model with 4 time steps requires only 5.67G Neuromorphic Synaptic Arithmetic Computation Effort (NS-ACE Shen et al. (2024), more details are shown in Appendix 6.4), a 94.6% reduction from the 104.32G NS-ACE of the full-precision model. This substantial decrease in NS-ACE indicates markedly improved energy efficiency of our model on neuromorphic hardware.
@@ 4.2.1 Impact of components of CIE on model accuracy
Table 4: Ablation study of CIE design on CIFAR-100. RF stands for reversible architecture, and IED stands for information enhancement distillation. †means that $X_{l}^{S,0}$ in Figure 3 is used for distillation and $X_{l}^{S,1}$ is used for classification. ‡means that $X_{l}^{S,0}$ is used for classification and $X_{l}^{S,1}$ is used for distillation.
@@
w/ RF & IED† 79.67 + 1.90
w/ RF & IED‡ 79.80 + 2.03

### v2

@@ Abstract
Transformer-based Spiking Neural Networks (SNNs) introduce a novel event-driven self-attention paradigm that combines the high performance of Transformers with the energy efficiency of SNNs. However, the larger model size and increased computational demands of the Transformer structure limit their practicality in resource-constrained scenarios. In this paper, we integrate binarization techniques into Transformer-based SNNs and propose the Binary Event-Driven Spiking Transformer, i.e. BESTformer. The proposed BESTformer can significantly reduce storage and computational demands by representing weights and attention maps with a mere 1-bit. However, BESTformer suffers from a severe performance drop from its full-precision counterpart due to the limited representation capability of binarization. To address this issue, we propose a Coupled Information Enhancement (CIE) method, which consists of a reversible framework and information enhancement distillation. By maximizing the mutual information between the binary model and its full-precision counterpart, the CIE method effectively mitigates the performance degradation of the BESTformer. Extensive experiments on static and neuromorphic datasets demonstrate that our method achieves superior performance to other binary SNNs, showcasing its potential as a compact yet high-performance model for resource-limited edge devices. The repository of this paper is available at https://github.com/CaoHLin/BESTFormer.
@@ 1 Introduction
Spiking Neural Networks (SNNs) have attracted significant attention as third-generation artificial neural networks, known for their high biological plausibility and low power consumption Maass (1997). The spiking neuron utilizes binary spikes as the fundamental units for information transmission and works in a sparse spike-driven manner Zhang et al. (2021). This sparse synaptic transmission in SNNs simplifies multiply-accumulate (MAC) operations into accumulate (AC) operations, thereby significantly enhancing computational efficiency Li et al. (2023); Xu et al. (2024a). Furthermore, the energy-efficient feature of SNNs has driven the development of neuromorphic hardware, such as TrueNorth Akopyan et al. (2015), Loihi Davies et al. (2018), and Tianjic Pei et al. (2019). However, despite the notable energy efficiency of SNNs, their performance in complex tasks still requires improvement.
Figure 1: Accuracy vs. NS-ACE & Model Size. Our method achieves superior computational and storage efficiency while outperforming other quantized SNNs on ImageNet. Neuromorphic Synaptic Arithmetic Computation Effort (NS-ACE) assesses SNN resource use in neuromorphic computing environments Shen et al. (2024).
In recent years, there have been several studies integrating Transformers into SNNs, leading to a series of high-performance models, such as Spikformer Zhou et al. (2023b), Spikingformer Zhou et al. (2023a), Spike-Driven Transformer v1 and v2 Yao et al. (2024a, b), and SNN-ViT Wang et al. (2025). Compared to convolutional architectures in SNNs, these Transformer-based models have demonstrated significant performance improvements Zhang et al. (2022). However, their advancements typically rely on large model size, which is accompanied by substantial memory storage and computational overhead, limiting their deployment on resource-constrained edge devices. Therefore, there is an urgent need for a compact yet high-performance Transformer-based SNN.
Quantization is a highly effective method for compressing large-scale models, which reduces model parameters from 32-bit to a low bit-width representation Wei et al. (2025). As an extreme form of quantization, binarization maximizes model size compression and accelerates computational speed by employing bitwise operations Qin et al. (2020a). Therefore, incorporating binarization with Transformer-based SNNs is promising for achieving an efficient and high-performance model. It is worth noting that current research on binarization in SNN domains is primarily focused on convolutional structures, while Transformer-based structures remain unexplored Yin et al. (2024); Wei et al. (2024); Liang et al. (2025).
@@ 2.1 Transformer-based SNNs
In recent years, there have been several studies integrating Transformer architectures into SNNs, leading to a series of high-performance SNN models. Spikeformer Li et al. (2022b) is the first to integrate the Transformer architecture with SNNs, however, it retains numerous floating-point operations, making it unsuitable for neuromorphic computation. Spikformer Zhou et al. (2023b) introduces the Spiking Self Attention (SSA) mechanism, which enhances both energy efficiency and performance of Transformer-based SNNs. Based on this, Spikingformer Zhou et al. (2023a) modifies the residual connection within it to achieve a purely spike-driven Vision Transformer, further enhancing model’s efficiency. Spike-driven Transformer Yao et al. (2024b) proposes a spike-driven self-attention mechanism with linear complexity, significantly reducing energy consumption. Then, to ensure versatility and high performance across various vision tasks, Yao et al. (2024a) expand the original architecture into Meta-SpikeFormer, also known as Spike-driven Transformer v2. Moreover, SNN-ViT Wang et al. (2025) achieves state-of-the-art performance in spiking vision tasks by introducing a saccadic self-attention mechanism specifically designed for spatio-temporal spike trains, maintaining linear computational complexity for edge applications. Despite much progress, these models are limited by substantial memory and computational overheads, underscoring the need for further compression to reach their full potential.
@@ 2.2 Quantization techniques in SNNs
Various approaches have been proposed to quantize SNNs to low-bits. Deng et al. (2021) employs spatiotemporal backpropagation (STBP) to directly train quantized SNNs and introduces the alternating direction method of multipliers (ADMM) to solve the performance degradation caused by quantization. Then, to further enhance the performance, Yoo and Jeong (2023) uses constrained backpropagation (CBP) with the Lagrangian function as an objective function to quantize SNNs in training. As an extreme form of quantization, binarization has also been widely studied. Qiao et al. (2021) presents a weight-binarized SNN to efficiently process event-based data, addressing the training demand of neuromorphic hardware for event data. Moreover, Pei et al. (2023) proposes the accuracy loss estimator and binary weight optimization to achieve ultra-low latency adaptive local binary SNNs, which reduce memory storage by over 20% while still maintaining high recognition accuracy. Recently, Wei et al. (2024) introduces a quantized SNN (Q-SNN) that reduces both weight and membrane potential representation, they also propose a weight-spike dual regulation (WS-DR) method to enhance the performance of Q-SNN. Despite the effectiveness of these binarization methods in SNNs, they mainly focus on spiking convolutional architectures, while Transformer-based SNNs have not been explored.
@@ 3 Method
Table 1: A simple example of binarizing Attn with $\operatorname{boolean}$ function, HR-LIF, and SR-LIF. The number of split patches NN, time step $T$, threshold $V_{th}$, and time constant $\tau$ are set to 1, 4, 1, 0.5, respectively.
@@ 3.1.1 Weight binarization
where $\tau$ is time constant factor, $\text{U}_{l}[t]$ is the membrane potential of neurons in layer $l$ at time t1t-1, $\tilde{\text{U}}_{l}[t]$ is its intermediate representation, and $\text{X}_{l}[t]\in\mathbb{R}^{C\times H\times W}$ is the input current. $\text{X}_{l}[t]$ is integrated by presynaptic neurons, described as:
@@
Neurons reset their membrane potential after emitting a spike. Typically, we set the reset potential to 0:
@@
where $\otimes$ is efficient bitwise operations, theoretically offering a $32\times$ memory saving and speedup compared with 32-bit operations Rastegari et al. (2016).
Figure 2: The ‘LIF-B-Conv-BN’ structure of BESTformer and representation capability of variables in the structure. Value set indicates the collection of all values present in a variable. Set size indicates the size of a value set.
@@ 3.1.2 Attention binarization
$\text{B}_{\text{Attn}}=\lambda\cdot\text{SR-LIF}(\text{Attn}),$ (9)
@@ 3.2 Challenge analysis
Previous research indicates that the poor performance of binarized neural networks is attributable to their low representational capability Liu et al. (2018); Guo et al. (2022, 2024). In our research, we experimentally analyzed the average representation capability of the full-precision network and BESTformer on ImageNet-1k, respectively. We use the value set size of feature maps as the measure of representation capability, which refers to the maximum number of distinct values in a feature map. The specific results are shown in Figure 2.
@@ 3.3 Coupled information enhancement BESTformer
Due to the constrained information representation capability of the binarized model, the ideal binary model should aim to retain the information representation of their full-precision counterparts as much as possible, thus the mutual information between the binarized and full-precision models’ representations should be maximized Li et al. (2022a); Xu et al. (2023, 2024b). Therefore, we propose a knowledge distillation framework, optimizing the mutual information $\mathcal{I}$ between the student model $\mathcal{S}$ and the full-precision teacher model $\mathcal{T}$, which can be formalized as:
@@
$\mathcal{I}(\text{X}_{n}^{\mathcal{S}};\text{X}_{m}^{\mathcal{T}})=\mathcal{H}(\text{X}_{n}^{\mathcal{S}})-\mathcal{H}(\text{X}_{n}^{\mathcal{S}}|\text{X}_{m}^{\mathcal{T}}),$ (11)
where $\mathcal{H}(\mathbf{X})=-\int p(\boldsymbol{x})\log p(\boldsymbol{x})d\boldsymbol{x}$ and $p(\boldsymbol{x})$ denotes the probability density function of the random variable. We employed coupled information enhancement methods to optimize the objective: (1) maximizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$ to its upper bound by reversible framework Gomez et al. (2017), and (2) minimizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}}|\text{X}_{m}^{\mathcal{T}})$ by information-enhanced distillation. The overall architecture diagram of applying the CIE method to BESTformer is shown in Figure 3.
@@ 3.3.1 Reversible framework
The information entropy of the model’s feature maps exhibits a decreasing trend as the number of layers increases. This can be shown explicitly by the inequality in Proposition 1.
@@
Proposition 1 shows that information retained by the model tends to decrease as the neural network deepens, which is contrary to our objective of maximizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$. To address this issue, we utilize a reversible forward mapping in the encoder, as the reversible connection shown in Figure 3. We define forward mapping $\Phi_{l}(\text{X}_{l-1}^{\mathcal{S},0},\text{X}_{l-1}^{\mathcal{S},1})=(\text{X}_{l}^{\mathcal{S},0},\text{X}_{l}^{\mathcal{S},1})$ as:
@@
Inverse mapping does not directly engage in network’s operations, but it indirectly confirms that the information entropy within the reversible framework remains constant across all encoder blocks. Therefore, $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$ reaches its upper bound $\mathcal{H}(\text{X}_{0}^{\mathcal{S}})$ as demonstrated by Proposition 2, which aligns with the objective of maximizing $\mathcal{H}(\text{X}_{n}^{\mathcal{S}})$. The detailed proofs of Proposition 1 and 2 are given in Supplementary Materials.
@@
Figure 4: Illustration of the forward and inverse process of the proposed reversible framework. The inverse process indicates that the inputs can be reconstructed from the outputs, i.e., this framework is reversible and no information is lost.
@@ 3.3.2 Information enhanced distillation
This dual-head design offers a significant advantage over traditional distillation methods by decoupling the gradient backpropagation for distillation and classification at the head level. This decoupling reduces mutual interference between the two gradients during backpropagation, facilitating more effective parameter updates. It is worth noting that $\text{X}_{n}^{\mathcal{S},1}$, being the output of a deeper network compared to $\text{X}_{n}^{\mathcal{S},0}$, is more susceptible to overfitting when used for classification. Therefore, we strategically feed $\text{X}_{n}^{\mathcal{S},1}$ into the distillation head. This approach leverages the teacher model’s output, which encapsulates ‘error information’, thereby alleviating the overfitting problem and enhancing model’s generalization capabilities.
@@ 4 Experiments
In this section, we first assess the classification performance of the proposed BESTformer with the CIE method on small-scale datasets, including CIFAR Krizhevsky et al. (2009), CIFAR10-DVS Li et al. (2017). Following this, we evaluate the method’s performance on large-scale image dataset, ImageNet-1K Deng et al. (2009) to verify the scalability of our approach. Finally, we perform a series of ablation studies to validate the effectiveness of our method. The implementation details are provided in Supplementary Materials.
@@ 4.1.2 Results on ImageNet-1k classification
Compared to full-precision models, our method significantly reduces model size and computational efficiency. According to Table 3, our 1-bit Bestformer-8-512 model is 5.57MB, approximately 10 times smaller than its full-precision counterpart (59.36MB). In terms of computational efficiency, our 1-bit Bestformer-8-512 model with 4 time steps requires only 5.67G Neuromorphic Synaptic Arithmetic Computation Effort (NS-ACE Shen et al. (2024), more details are shown in Supplementary Materials), a 94.6% reduction from the 104.32G NS-ACE of the full-precision model. This substantial decrease in NS-ACE indicates markedly improved energy efficiency of our model on neuromorphic hardware.
@@ 4.2 Ablation study
Table 3: Comparison of resource consumption on ImageNet-1k. $\dagger$ The results for the full-precision model are taken from Zhou et al. (2023a). $\ddagger$ The quantized results are taken from Shen et al. (2024).
@@ 4.2.1 Impact of components of CIE on model accuracy
[…]
Table 4: Ablation study of CIE design on CIFAR-100. RF stands for reversible architecture, and IED stands for information enhancement distillation. Symbol $\dagger$ means that $X_{l}^{S,0}$ in Figure 3 is used for distillation and $X_{l}^{S,1}$ is used for classification. Symbol $\ddagger$ means that $X_{l}^{S,0}$ is used for classification and $X_{l}^{S,1}$ is used for distillation.